Q: solve for x and y : (a+b)x + (a-b)y=2ab (a+b)x - (a-b)y =ab.

Grade:10

1 Answers

Rinkoo Gupta
9 years ago
(a+b)x+(a-b)y=ab
2ab(a+b)x-(a-b)y=ab
Adding both eqs, we get
(1+2ab)(a+b)x=2ab
x=2ab/(2ab+1)(a+b)
putting the value of x in eq (1)
(a+b).2ab/(2ab+1)(a+b) +(a-b)y=ab
(a-b)y=ab-2ab/(2ab+1)
=ab(2ab-1)/(2ab+1)
y=ab(2ab-1)/(2ab+1)(a-b)

Thanks & Regards
Rinkoo Gupta
AskIITians Faculty